# captionUnknown document class (or package)

### v1

@@ VI-B UGV Experiments
Trials: A full 22 datasets were collected with all three agents, each with 6 antennas, moving in various ways within the mocap space. Ranging measurements are performed by each pair of antennas between agents at 25Hz. Together the datasets total to nearly 6h (effectively creating over 200h of pairwise measurements). This dataset was used for Figures 1(a) and 1(b). Tables IV and IV show results for the 5 trials where all three agents are continuously moving in various arbitrary patterns with max positional and angular velocities of 1m/s and 1rad/s respectively.
Parameters: Our approach, outlined in Section V and culminating in Equation 9 and Algorithm 1, is configured as such: (1) Our mean bias correction function, $\bar{d}^{A_{i}}_{B_{j}}(\mathbf{T}^{A}_{B})$, is set to the learned 6-degree polynomial shown in Figure 1(b). This incorporates the observed mean bias with respect to elevation into our approach. (2) The current range measurements, $\tilde{d}^{A_{i}}_{B_{j}}$, and pose estimates are smoothed by a moving average filter with a 1s and 4s window length respectively. (3) Our loss function, $\ell$, is selected to be the Huber loss $\rho_{\delta}(a)$ with $\delta=0.06$. As with [20, 21, 17], Huber loss was selected due to its reduced outliers sensitivity compared to a traditional squared error loss [34]. The parameter $\delta=0.06$ was selected after doing a parameter search over our results.
Discussion of Results: Results are summarized for Absolute Position Error (APE) and Absolute Heading Error (AHE) in Figures 4(a) and 4(b) and Tables IV and IV. To demonstrate the value of individual algorithmic decisions, the tables toggles el_bias, z_fixed, Huber, where the final row represents the proposed approach. Specifically, a check indicates:
@@
The tables show the proposed approach consistently improves the mean APE by an average of 19% over the second best APE approach (i.e., without the el_bias, orange), while constraining $z$ leads to the largest APE gains. Unlike APE, there is not a clear best AHE approach. That being said, the proposed approach is only at most $1^{\circ}$ behind the best mean AHE approach for any given trial.
VII Conclusion
This work presents a UWB-ranging inter-agent relative pose estimation system that minimizes communication load. Although challenging to do a direct comparison, other state-of-the-art approaches achieve APE and AHE on the order of 0.10m and $1^{\circ}$ respectively (see Section II), but rely heavily on continuously transmitted measurements which leading to potential scaling issues with increased team size and/or decreased communication network capability By comparison, this work achieves competitive 3D results, averaged across all trials a mean APE and AHE of 0.24m and $9.5^{\circ}$ respectively, while requiring no continuously transmitted measurements. Our work outperforms the most-similar recent work [23], in part due to our learned measurement mean bias error, $\bar{d}^{A_{i}}_{B_{j}}(\mathbf{T}^{A}_{B})$, which greatly improves our APE by a average of 19% across our experimental trials. In the future, this mean bias error estimate can be further improved as a neural network – that being said, a key takeaway is that most UWB systems would benefit from a pose-dependent UWB ranging error model. Furthermore, future work will include additional flights on multiple larger drones in expansive outdoor environments and integration of this system into a larger resource-aware distributed SLAM pipeline.

### v3

@@ V-B UGV Experiments
Trials: A full 22 datasets were collected with all three agents, each with 6 antennas, moving in various ways within the mocap space. Ranging measurements are performed by each pair of antennas between agents at 25Hz. Together the datasets total to nearly 6h (effectively creating over 200h of pairwise measurements). These datasets were used for Figures 2(a) and 2(b). Tables V and V show results for the 5 trials where all three agents are continuously moving in various arbitrary patterns with max positional and angular velocities of 1m/s and 1rad/s respectively. Additionally, two outdoor trials were conducted without mocap to verify the system generalizes to non-laboratory spaces and can operate for extended periods of time.
Parameters: (1) Equation 9 is solved by trust-constr as provided by scipy’s minimize function [25]. [^9] As in [10], initialized with previous estimate if available, otherwise $\langle\texttt{avg}(\tilde{\mathbf{d}}^{A}_{B}),0,0\rangle$. Here $\texttt{avg}(\tilde{\mathbf{d}}^{A}_{B})$ is the mean of the current measurements. State estimation is run at a fixed 1Hz. [^10] Each robot’s onboard i7-4710HQ processor can run estimation at ~5Hz serially and ¿25Hz in parallel (i.e., faster than the UWB sensor rate). (2) Our mean bias correction, $\bar{d}^{A_{i}}_{B_{j}}(\mathbf{T}^{A}_{B})$, is the learned 6-degree polynomial with respect to elevation shown in Figure 2(b). (3) The current range measurements, $\tilde{d}^{A_{i}}_{B_{j}}$, and pose estimates are smoothed by a 1s and 4s moving average filter respectively. (4) Our loss function, $\ell$, is selected to be the Huber loss $\rho_{\delta}(a)$ with $\delta=0.06$. As with [12, 13, 9], Huber loss was selected due to its reduced outlier sensitivity [26].
Discussion of Results: Results are shown in Tables V and V, Figure 5, and Figure 6. To demonstrate the value of individual algorithmic decisions, the tables toggle el_bias, z_fixed, Huber, where the final row (red) represents the proposed approach. Specifically, a check indicates:
@@
Our data shows the proposed approach (red) provides a 9$\times$ improvement over a direct NLLS trilateration (purple), similar to that used by [15]. The addition of the el_bias improves our approach’s mean APE by an average of 19% (i.e., red vs orange). Overall, constraining $z$ leads to the largest APE gains (as predicted by Table II). Unlike APE, there is not a clear best approach for AHE, but the proposed (red) is only at most $1^{\circ}$ behind the best for any given trial. Additionally, Figure 6 justifies our minimal communication protocol, demonstrating that even when operating with maximal altitude error within the experimentally demonstrated error tolerances (i.e., ±$\pm 0.10$m from Figure 4(a)), red outperforms other formulations. Finally, the two outdoor trials, highlighted in the video, demonstrated comparable position accuracy to the laboratory trials as well as the long-term stability and reliability of the system.
VI Conclusion
This work presents a UWB-ranging inter-agent relative pose estimation system that minimizes communication load. Our work outperforms the most-similar recent work [15] (APEs between 0.40m and 1m). We also remain competitive with other state-of-the-art approaches that have significantly higher communication costs (APE and AHE on the order of 0.10m and $1^{\circ}$ respectively), whose heavy reliance on continuously transmitted measurements could lead to scaling issues with increased team size and/or decreased communication network capability.
By comparison, this work achieves a mean APE and AHE of 0.24m and $9.5^{\circ}$ respectively across our experimental trials, nearly a 9$\times$ improvement over a direct NLLS trilateration approach while requiring no continuously transmitted measurements. Our learned measurement mean bias error correction, $\bar{d}^{A_{i}}_{B_{j}}(\mathbf{T}^{A}_{B})$, greatly improves our APE by an average of 19% across our experimental trials. In the future, this bias correction can be further improved as a neural network, potentially dependent on more than just relative elevation – that being said, a key takeaway is that most UWB systems would benefit from a pose-dependent UWB ranging error model. Furthermore, future work will include additional flights on multiple larger drones in expansive outdoor environments, integration of this system into a larger resource-aware distributed SLAM pipeline, and locally synchronizing UWB hardware as in [27].